# Effects of personality steering on cooperative behavior in Large Language Model agents

### v1

@@ Abstract
Large language models (LLMs) are increasingly used as autonomous agents in strategic and social interactions. Although recent studies suggest that assigning personality traits to LLMs can influence their behavior, how personality steering affects cooperation under controlled conditions remains unclear. In this study, we examine the effects of personality steering on cooperative behavior in LLM agents using repeated Prisoner’s Dilemma games. Based on the Big Five framework, we first measure basic personality profiles of three models, GPT-3.5-turbo, GPT-4o, and GPT-5, using the Big Five Inventory. We then compare behavior under baseline and personality-informed conditions, and further analyze the effects of independently manipulating each personality dimension to extreme values. Our results show that agreeableness is the dominant factor promoting cooperation across all models, while other personality traits have limited impact. Explicit personality information increases cooperation but can also raise vulnerability to exploitation, particularly in earlier-generation models. In contrast, later-generation models exhibit more selective cooperation. These findings indicate that personality steering acts as a behavioral bias rather than a deterministic control mechanism.
@@ 1 Introduction
Multi-agent systems have long been studied and utilized to make decisions on behalf of humans. Traditional multi-agent systems were designed to achieve rational or cooperative behavior under specific environments, based on limited information-processing capabilities and predefined rules. Typical applications include price negotiations and contract adjustments, where multiple agents seek agreement while pursuing conflicting interests [1]; resource allocation scenarios, where competition and cooperation coexist over limited resources [2]; and marketplaces or auctions, where seller and buyer agents strategically propose prices [3, 1]. In recent years, the emergence of Large Language Models (LLMs) has substantially expanded the decision-making capabilities of agents. By employing LLMs as agents, flexible interpretation of information and reasoning through natural language become possible, enabling decision-making even in complex and context-dependent social situations that were previously difficult to handle. As a result, LLM-based multi-agent systems have demonstrated new potential in social interactions involving multiple stakeholders and in settings that require human-like judgment [4, 5, 6, 7].
@@
However, this shift in agent capabilities has also introduced new challenges. Interactions among LLM-based agents can be unpredictable [4] and may involve the risk of escalating unintended conflicts [8]. In strategic environments, optimal decision-making requires not only the maximization of immediate rewards but also careful consideration of relationships with other agents and the formation of long-term cooperative relationships. Such instability highlights the need for new approaches to guide and control the behavior and decision-making of LLMs in strategic settings. As a potential solution to this challenge, the psychological framework of personality has attracted increasing attention. In personality psychology, various theories have been proposed to explain individual behavioral patterns, among which the Big Five Personality Traits [9] are widely accepted as a theoretical framework that describes personality along five major dimensions: (1) Extraversion, which reflects sociability, activity level, and the tendency to seek stimulation; (2) Agreeableness, which captures empathy, cooperativeness, and tolerance toward others; (3) Conscientiousness, which represents responsibility, organization, and motivation toward goal achievement; (4) Neuroticism, which reflects emotional instability and sensitivity to stress; and (5) Openness to Experience, which captures openness to new experiences and ideas, as well as creativity. In human psychological research, individuals with high agreeableness tend to exhibit stronger cooperative behavior, while also being more vulnerable to exploitation [10]. Moreover, in strategic situations such as repeated Prisoner’s Dilemma games, individuals with higher agreeableness are more likely to cooperate in the early stages of interaction [11].
@@
At the same time, because the Big Five framework provides well-established quantitative measurement methods, it is also well suited as a basis for analyzing personality traits in LLMs, suggesting that such traits may influence decision-making in these models. Indeed, a previous study has reported that increasing traits such as agreeableness and conscientiousness in LLM agents promotes cooperative behavior, while simultaneously increasing vulnerability to exploitation by other agents, leading to lower overall payoffs-patterns that closely resemble those observed in humans [12]. However, this prior study has several important limitations. First, it typically begins by assigning or manipulating personality traits in LLMs without first quantitatively measuring the basic personality profiles inherently exhibited by the models, making it difficult to assess how personality interventions alter behavior relative to a model’s original tendencies. Second, because personality traits are often not specified in quantitative terms, the strength of personality manipulations cannot be systematically compared, which limits reproducibility and hinders meaningful comparisons across studies.
@@
To address these limitations, we investigate the relationship between personality traits and cooperative behavior in LLM agents under quantitatively controlled conditions using the Big Five Personality Traits framework. Specifically, we examine three objectives. First, we quantitatively measure the basic personality profiles inherently exhibited by different LLMs based on the Big Five framework, thereby clarifying model-specific characteristics and differences across model generations. Second, we examine how LLM agents behave in strategic settings such as repeated Prisoner’s Dilemma games when their measured personality traits are explicitly provided through prompts. Third, we analyze how cooperative behavior changes when each personality dimension is independently manipulated to extreme values while holding the remaining dimensions constant. Through these analyses, we aim to clarify how personality steering influences cooperative behavior in LLM agents and to provide a controlled framework for understanding the role of personality traits in LLM decision-making.
@@ 2 Methods
In this study, we investigate the relationship between personality traits and cooperative behavior in LLM agents using three widely used models developed by OpenAI: GPT-3.5-turbo, GPT-4o, and GPT-5. We conducted the experiments in three stages. First, we quantitatively measured the basic personality profiles of each model based on the Big Five Personality Traits using the Big Five Inventory (BFI-44) (Experiment 1). Second, we examined how LLM agents behave in strategic settings by having them play repeated Prisoner’s Dilemma games under two conditions: a baseline condition without explicit personality information and a condition in which the measured personality traits obtained in Experiment 1 were explicitly provided via prompts (Experiment 2). Third, we analyzed the effects of personality manipulation by independently setting each Big Five trait to extreme values while keeping the remaining traits constant, in order to assess which personality dimensions most strongly influenced cooperative behavior (Experiment 3).
@@ 2.1 Experiment 1: Measuring basic personality traits using the Big Five
In Experiment 1, we first measure the personality traits of each model based on the Big Five Personality Traits framework [9]. To quantitatively assess these personality traits in LLMs, we employed the Big Five Inventory (BFI), one of the most widely used psychometric instruments for measuring personality traits [9]. The BFI is a self-report questionnaire developed based on the Big Five theory and has been validated across diverse cultures and languages, demonstrating high reliability and validity. In this study, we used the BFI-44, a shortened version consisting of 44 items, with 8–10 items assigned to each personality dimension. Each item is rated on a five-point Likert scale ranging from 1 (“strongly disagree”) to 5 (“strongly agree”).
For each model, the 44 questionnaire items were presented sequentially, with instructions prompting the model to respond to each item using a numerical value from 1 to 5 (see Appendix A for the full prompt). We adopt the BFI prompt developed by Jiang et al. [13]. If a response did not conform to the required format (e.g., contains non-numeric characters), the question was presented again until a valid numerical response was obtained. After collecting responses to all items, scores for each personality dimension were computed following the standard BFI scoring procedure. Because LLM outputs involve stochasticity, we repeated the personality measurement five times for each model and used the average scores across the five runs to improve score stability. Through this procedure, we obtained quantitative and reproducible basic personality profiles for each LLM.
@@ 2.2 Experiment 2: Behavioral differences with explicit personality information
In Experiment 2, we examine how LLM agents behave in strategic environments by having them play repeated Prisoner’s Dilemma (RPD) games. We considered two experimental conditions. In the baseline condition, no explicit personality information was provided to the LLMs. In the personality-informed condition, the personality traits measured in Experiment 1 were explicitly provided to the LLMs via prompts. By comparing these two conditions, we assessed whether and how explicitly supplied personality information influenced cooperative behavior.
@@
The Prisoner’s Dilemma is a canonical game-theoretic model that captures social dilemmas arising from the tension between cooperation and defection. In each round, two players simultaneously choose either to cooperate (C) or defect (D). Payoffs are determined by the joint action of both players according to a payoff matrix satisfying $T>R>P>S$ , where $T$ denotes the temptation to defect, $R$ the reward for mutual cooperation, $P$ the punishment for mutual defection, and $S$ the payoff received by a cooperator when the opponent defects. In this study, we used the standard payoff values $T=5$, $R=3$, $P=1$, and $S=0$.
@@
In the RPD game, the same pair of players interacts repeatedly, and it is assumed that each player can observe and remember past actions and outcomes. This repeated structure allows players to condition their current decisions on previous interactions, thereby enabling the emergence of cooperative behavior. Both theoretical [14, 15, 16, 17, 18, 19, 20] and experimental [21, 22, 23, 24, 25] studies have demonstrated that cooperation can be sustained in repeated Prisoner’s Dilemma games under appropriate conditions.
@@
We first conducted the RPD experiments under the baseline condition to characterize the intrinsic cooperative tendencies of each LLM. In this condition, the prompt included only the rules and objectives of the RPD game and did not contain any information related to personality traits. The prompt was constructed based on that proposed by Fontana et al. [26]. Each LLM agent played 10 independent trials, each consisting of 10 rounds of the RPD game. Trials were independent, and outcomes from previous trials did not affect subsequent trials.
@@
In each trial, the LLM agent played against a set of fixed opponent strategies that were hard-coded and commonly used in studies of the RPD game. These strategies include: Always Cooperate (ALLC), which selects cooperation in every round; Always Defect (ALLD), which selects defection in every round; Random (RANDOM), which randomly chooses between cooperation and defection in each round; Tit-for-Tat (TFT), which cooperates in the first round and then replicates the opponent’s previous action; and Grim Trigger (GRIM), which cooperates initially but switches permanently to defection after the opponent defects once. Together, these strategies cover a broad range of cooperative and non-cooperative behaviors.
@@
In the personality-informed condition, we repeated the same RPD experiments while explicitly providing each LLM with its measured Big Five personality scores obtained in Experiment 1. The prompt specifies the personality traits in the form: “Your personality traits are as follows: Openness (O), Conscientiousness (C), Extraversion (E), Agreeableness (A), and Neuroticism (N),” followed by the corresponding numerical scores (see Appendix for prompt details). Under this condition, the LLM was instructed to make decisions while being aware of its own personality traits.
@@
For each experimental condition, we computed two evaluation metrics. The average cooperation rate is defined as the proportion of cooperative choices across all rounds, indicating the degree of cooperativeness exhibited by the LLM. The average cumulative payoff is defined as the mean total payoff accumulated across all rounds, reflecting the overall performance of the LLM. By comparing these metrics between the baseline and personality-informed conditions, we evaluated the behavioral impact of explicitly providing personality information to LLM agents.
@@ 2.3 Experiment 3: Effects of extreme personality manipulation
In each experimental condition, only one personality trait was manipulated. For example, when manipulating agreeableness, the agreeableness score is set to either 1 (minimum) or 5 (maximum), while the scores for openness, conscientiousness, extraversion, and neuroticism remain unchanged from the values measured in Experiment 1. By applying this manipulation independently to each of the five personality dimensions and considering both extreme values, we obtain a total of ten experimental conditions (five dimensions $\times$ two values).
For each condition, the modified personality scores are explicitly provided to the LLM agents via prompts in the same format as in the personality-informed condition of Experiment 2. The agents then play the RPD games under identical settings, including the same opponent strategies, number of rounds, and number of trials. Using the same evaluation metrics as in Experiment 2, the average cooperation rate and the average cumulative payoff, we assessed how extreme manipulation of individual personality traits influenced cooperative behavior.
@@ 2.4 LLM Settings
In this study, we used three large language models developed by OpenAI: GPT-3.5-turbo, GPT-4o, and GPT-5. These models represent successive generations of LLMs, with larger model numbers corresponding to more recent releases.
For GPT-3.5-turbo and GPT-4o, we set the temperature parameter to 0.7. The temperature controls the degree of randomness in model outputs, with lower values producing more deterministic responses and higher values increasing stochasticity. The value of 0.7 is chosen to balance response diversity and consistency, and is consistent with the setting used in prior work [26].
For GPT-5, we conducted experiments with the reasoning effort set to minimal and the verbosity set to low. The reasoning effort parameter controls the amount of computational effort allocated to the model’s internal reasoning process; setting it to minimal simplifies the reasoning process and improves response speed. The verbosity parameter controls the level of detail in the model’s output; setting it to low ensures that only the minimum necessary information is produced.
@@ 3.1 BFI scores
We first report the results of Experiment 1, in which the personality traits of each LLM were measured using the Big Five Inventory. Table 1 presents the mean scores and standard deviations obtained from five repeated measurements for each model. For comparison, the table also includes human data reported by Srivastava et al. [27], where the original POMP scores were rescaled to the same 1–5 scale used in this study.
@@
Openness (O) 4.60 (0.14) 4.78 (0.04) 4.78 (0.13) 3.98 (0.66)
Conscientiousness (C) 4.09 (0.05) 4.09 (0.09) 4.78 (0.21) 3.55 (0.73)
Extraversion (E) 3.80 (0.19) 3.00 (0.29) 3.10 (0.27) 3.18 (0.90)
Agreeableness (A) 4.15 (0.19) 4.31 (0.31) 4.24 (0.21) 3.66 (0.72)
Neuroticism (N) 1.88 (0.13) 1.89 (0.29) 2.03 (0.10) 3.04 (0.88)
@@
Across all models, the five measurements yielded highly consistent scores, with relatively small standard deviations. Specifically, the standard deviations ranged from 0.05 to 0.19 for GPT-3.5-turbo, from 0.04 to 0.31 for GPT-4o, and from 0.10 to 0.27 for GPT-5. These values are substantially smaller than those observed in humans (0.66–0.90), indicating that each LLM exhibits a stable and internally consistent personality profile. This consistency suggests that the measured personality traits reliably characterize each model.
In contrast, the human scores cluster around the mid-range of the scale, with larger standard deviations, reflecting substantial inter-individual variability in human personality traits. Compared to human averages, all LLMs exhibit higher scores in openness, conscientiousness, and agreeableness, as well as lower scores in neuroticism. This pattern suggests that LLMs tend to exhibit personality profiles that are systematically shifted toward higher scores in certain traits compared to humans, potentially reflecting biases toward socially desirable response patterns present in their training data.
Next, we examined the patterns observed across individual personality dimensions. Overall, all models exhibited broadly similar tendencies, characterized by relatively high scores in openness, conscientiousness, and agreeableness, and low scores in neuroticism. This consistency across models suggests that contemporary LLMs share common personality profiles as measured by the BFI framework. A particularly notable result is observed for conscientiousness. While GPT-3.5-turbo and GPT-4o show identical mean scores for conscientiousness (4.09), GPT-5 exhibits a substantially higher score (4.78). One possible explanation for this difference is the technical improvements reported for GPT-5, including a significant reduction in hallucinated responses and enhanced reasoning consistency compared to earlier models [28]. Such improvements may contribute to more reliable, goal-oriented, and internally consistent responses, which are reflected in higher conscientiousness scores.

### v2

@@ Abstract
Large language models (LLMs) are increasingly used as autonomous agents in strategic and social interactions. Although recent studies suggest that assigning personality traits to LLMs can influence their behavior, how personality steering affects cooperation under controlled conditions remains unclear. In this study, we examine the effects of personality steering on cooperative behavior in LLM agents using repeated Prisoner’s Dilemma games. Based on the Big Five framework, we first measure basic personality scores of three models, GPT-3.5-turbo, GPT-4o, and GPT-5, using the Big Five Inventory. We then compare behavior under baseline and personality-informed conditions, and further analyze the effects of independently manipulating each personality dimension to extreme values. Our results show that agreeableness is the dominant factor promoting cooperation across all models, while other personality traits have limited impact. Explicit personality information increases cooperation but can also raise vulnerability to exploitation, particularly in earlier-generation models. In contrast, later-generation models exhibit more selective cooperation. These findings indicate that personality steering acts as a behavioral bias rather than a deterministic control mechanism.
@@ 1 Introduction
Multi-agent systems have long been studied and utilized to make decisions on behalf of humans. Traditional multi-agent systems were designed to achieve rational or cooperative behavior under specific environments, based on limited information-processing capabilities and predefined rules. Typical applications include price negotiations and contract adjustments, where multiple agents seek agreement while pursuing conflicting interests [Shoham2009]; resource allocation scenarios, where competition and cooperation coexist over limited resources [Perolat2017]; and marketplaces or auctions, where seller and buyer agents strategically propose prices [Wellman2016, Shoham2009]. In recent years, the emergence of Large Language Models (LLMs) has substantially expanded the decision-making capabilities of agents. By employing LLMs as agents, flexible interpretation of information and reasoning through natural language become possible, enabling decision-making even in complex and context-dependent social situations that were previously difficult to handle. As a result, LLM-based multi-agent systems have demonstrated new potential in social interactions involving multiple stakeholders and in settings that require human-like judgment [Park2023, Newsham2025, Zeng2025, Akata2025].
@@
However, this shift in agent capabilities has also introduced new challenges. Interactions among LLM-based agents can be unpredictable [Park2023] and may involve the risk of escalating unintended conflicts [Rivera2024]. In strategic environments, optimal decision-making requires not only the maximization of immediate rewards but also careful consideration of relationships with other agents and the formation of long-term cooperative relationships. Such instability highlights the need for new approaches to guide and control the behavior and decision-making of LLMs in strategic settings. As a potential solution to this challenge, the psychological framework of personality has attracted increasing attention. In personality psychology, various theories have been proposed to explain individual behavioral patterns, among which the Big Five Personality Traits [John1991] are widely accepted as a theoretical framework that describes personality along five major dimensions: (1) Extraversion, which reflects sociability, activity level, and the tendency to seek stimulation; (2) Agreeableness, which captures empathy, cooperativeness, and tolerance toward others; (3) Conscientiousness, which represents responsibility, organization, and motivation toward goal achievement; (4) Neuroticism, which reflects emotional instability and sensitivity to stress; and (5) Openness to Experience, which captures openness to new experiences and ideas, as well as creativity. In human psychological research, individuals with high agreeableness tend to exhibit stronger cooperative behavior, while also being more vulnerable to exploitation [Thielmann2020]. Moreover, in strategic situations such as repeated Prisoner’s Dilemma games, individuals with higher agreeableness are more likely to cooperate in the early stages of interaction [Kagel2014].
@@
At the same time, because the Big Five framework provides well-established quantitative measurement methods, it is also well suited as a basis for analyzing personality traits in LLMs, suggesting that such traits may influence decision-making in these models. Indeed, a previous study has reported that increasing traits such as agreeableness and conscientiousness in LLM agents promotes cooperative behavior, while simultaneously increasing vulnerability to exploitation by other agents, leading to lower overall payoffs-patterns that closely resemble those observed in humans [Ong2025]. However, this prior study has several important limitations. First, it typically begins by assigning or manipulating personality traits in LLMs without first quantitatively measuring the basic personality scores inherently exhibited by the models, making it difficult to assess how personality interventions alter behavior relative to a model’s original tendencies. Second, because personality traits are often not specified in quantitative terms, the strength of personality manipulations cannot be systematically compared, which limits reproducibility and hinders meaningful comparisons across studies.
@@
To address these limitations, we investigate the relationship between personality traits and cooperative behavior in LLM agents under quantitatively controlled conditions using the Big Five Personality Traits framework. Specifically, we examine three objectives. First, we quantitatively measure the basic personality scores inherently exhibited by different LLMs based on the Big Five framework, thereby clarifying model-specific characteristics and differences across model generations. Second, we examine how LLM agents behave in strategic settings such as repeated Prisoner’s Dilemma games when their measured personality traits are explicitly provided through prompts. Third, we analyze how cooperative behavior changes when each personality dimension is independently manipulated to extreme values while holding the remaining dimensions constant. Through these analyses, we aim to clarify how personality steering influences cooperative behavior in LLM agents and to provide a controlled framework for understanding the role of personality traits in LLM decision-making.
@@ 2 Methods
In this study, we investigate the relationship between personality traits and cooperative behavior in LLM agents using three widely used models developed by OpenAI: GPT-3.5-turbo, GPT-4o, and GPT-5. We conduct the experiments in three stages. First, we quantitatively measure the basic personality scores of each model based on the Big Five Personality Traits using the Big Five Inventory (BFI-44) (Experiment 1). Second, we examine how LLM agents behave in strategic settings by having them play repeated Prisoner’s Dilemma games under two conditions: a baseline condition without explicit personality information and a condition in which the measured personality traits obtained in Experiment 1 are explicitly provided via prompts (Experiment 2). Third, we analyze the effects of personality manipulation by independently setting each Big Five trait to extreme values while keeping the remaining traits constant, in order to assess which personality dimensions most strongly influence cooperative behavior (Experiment 3).
@@ 2.1 Experiment 1: Measuring basic personality traits using the Big Five
In Experiment 1, we first measure the personality traits of each model based on the Big Five Personality Traits framework [John1991]. To quantitatively assess these personality traits in LLMs, we employ the Big Five Inventory (BFI), one of the most widely used psychometric instruments for measuring personality traits [John1991]. The BFI is a self-report questionnaire developed based on the Big Five theory and has been validated across diverse cultures and languages, demonstrating high reliability and validity. In this study, we use the BFI-44, a shortened version consisting of 44 items, with 8–10 items assigned to each personality dimension. Each item is rated on a five-point Likert scale ranging from 1 (“strongly disagree”) to 5 (“strongly agree”).
For each model, the 44 questionnaire items are presented sequentially, with instructions prompting the model to respond to each item using a numerical value from 1 to 5 (see Appendix A for the full prompt). We adopt the BFI prompt developed by Jiang et al. [Jiang2024]. If a response does not conform to the required format (e.g., contains non-numeric characters), the question is presented again until a valid numerical response is obtained. After collecting responses to all items, scores for each personality dimension are computed following the standard BFI scoring procedure. Because LLM outputs involve stochasticity, we repeat the personality measurement 20 times for each model and use the average scores across the 20 runs to improve score stability. Through this procedure, we obtain quantitative and reproducible basic personality scores for each LLM.
@@ 2.2 Experiment 2: Behavioral differences with explicit personality information
In Experiment 2, we examine how LLM agents behave in strategic environments by having them play repeated Prisoner’s Dilemma (RPD) games. We consider two experimental conditions. In the baseline condition, no explicit personality information is provided to the LLMs. In the personality-informed condition, the personality traits measured in Experiment 1 are explicitly provided to the LLMs via prompts. By comparing these two conditions, we assess whether and how explicitly supplied personality information influence cooperative behavior.
@@
The Prisoner’s Dilemma is a canonical game-theoretic model that captures social dilemmas arising from the tension between cooperation and defection. In each round, two players simultaneously choose either to cooperate (C) or defect (D). Payoffs are determined by the joint action of both players according to a payoff matrix satisfying $T>R>P>S$ , where $T$ denotes the temptation to defect, $R$ the reward for mutual cooperation, $P$ the punishment for mutual defection, and $S$ the payoff received by a cooperator when the opponent defects. In this study, we use the standard payoff values $T=5$, $R=3$, $P=1$, and $S=0$.
@@
In the RPD game, the same pair of players interacts repeatedly, and it is assumed that each player can observe and remember past actions and outcomes. This repeated structure allows players to condition their current decisions on previous interactions, thereby enabling the emergence of cooperative behavior. Both theoretical [Friedman1971, Kreps1982, Axelrod1981, Axelrod1984, Nowak1992, Nowak1993, Kandori1992] and experimental [Selten1986, Andreoni1993, Bo2005, Camera2009, Wang2017] studies have demonstrated that cooperation can be sustained in the RPD games under appropriate conditions.
@@
We first conduct the RPD experiments under the baseline condition to characterize the intrinsic cooperative tendencies of each LLM. In this condition, the prompt include only the rules and objectives of the RPD game and does not contain any information related to personality traits. The prompt is constructed based on that proposed by Fontana et al. [Fontana2024]. The full prompt details are provided in Appendix B. Each LLM agent play 100 independent trials, each consisting of 10 rounds of the RPD game. Trials are independent, and outcomes from previous trials do not affect subsequent trials.
@@
In each trial, the LLM agent play against a set of fixed opponent strategies that are hard-coded and commonly used in studies of the RPD game. These strategies include: Always Cooperate (ALLC), which selects cooperation in every round; Always Defect (ALLD), which selects defection in every round; Random (RANDOM), which randomly chooses between cooperation and defection in each round; Tit-for-Tat (TFT), which cooperates in the first round and then replicates the opponent’s previous action; and Grim Trigger (GRIM), which cooperates initially but switches permanently to defection after the opponent defects once. Together, these strategies cover a broad range of cooperative and non-cooperative behaviors.
@@
In the personality-informed condition, we repeat the same RPD experiments while explicitly providing each LLM with its measured Big Five personality scores obtained in Experiment 1. The prompt specifies the personality traits in the form: “Your personality traits are as follows: Openness (O), Conscientiousness (C), Extraversion (E), Agreeableness (A), and Neuroticism (N),” followed by the corresponding numerical scores (see Appendix C for prompt details). Under this condition, the LLM is instructed to make decisions while being aware of its own personality traits.
@@
For each experimental condition, we compute two evaluation metrics. The average cooperation rate is defined as the proportion of cooperative choices across all rounds, indicating the degree of cooperativeness exhibited by the LLM. The average cumulative payoff is defined as the mean total payoff accumulated across all rounds, reflecting the overall performance of the LLM. By comparing these metrics between the baseline and personality-informed conditions, we evaluate the behavioral impact of explicitly providing personality information to LLM agents.
@@ 2.3 Experiment 3: Effects of extreme personality manipulation
In each experimental condition, only one personality trait is manipulated. For example, when manipulating agreeableness, the agreeableness score is set to either 1 (minimum) or 5 (maximum), while the scores for openness, conscientiousness, extraversion, and neuroticism remain unchanged from the values measured in Experiment 1. By applying this manipulation independently to each of the five personality dimensions and considering both extreme values, we obtain a total of ten experimental conditions (five dimensions $\times$ two values).
For each condition, the modified personality scores are explicitly provided to the LLM agents via prompts in the same format as in the personality-informed condition of Experiment 2. The agents then play the RPD games under identical settings, including the same opponent strategies, number of rounds, and number of trials. Using the same evaluation metrics as in Experiment 2, the average cooperation rate and the average cumulative payoff, we assess how extreme manipulation of individual personality traits influence cooperative behavior.
@@ 2.4 LLM settings
In this study, we use three large language models developed by OpenAI: GPT-3.5-turbo, GPT-4o, and GPT-5. These models represent successive generations of LLMs, with larger model numbers corresponding to more recent releases.
For GPT-3.5-turbo and GPT-4o, we set the temperature parameter to 0.7. The temperature controls the degree of randomness in model outputs, with lower values producing more deterministic responses and higher values increasing stochasticity. The value of 0.7 is chosen to balance response diversity and consistency, and is consistent with the setting used in prior work [Fontana2024].
For GPT-5, we conduct experiments with the reasoning effort set to minimal and the verbosity set to low. The reasoning effort parameter controls the amount of computational effort allocated to the model’s internal reasoning process; setting it to minimal simplifies the reasoning process and improves response speed. The verbosity parameter controls the level of detail in the model’s output; setting it to low ensures that only the minimum necessary information is produced.
@@ 3.1 BFI scores
We first report the results of Experiment 1, in which the personality traits of each LLM were measured using the Big Five Inventory. Table 1 presents the mean scores and standard deviations obtained from 20 repeated measurements for each model. For comparison, the table also includes human data reported by Srivastava et al. [Srivastava2003], where the original POMP scores were rescaled to the same 1–5 scale used in this study. Figure 1 provides a radar-plot visualization of the average Big Five personality scores shown in Table 1.
@@
Openness (O) 4.58 (0.12) 4.68 (0.25) 4.69 (0.17) 3.98 (0.66)
Conscientiousness (C) 4.06 (0.12) 4.12 (0.32) 4.69 (0.16) 3.55 (0.73)
Extraversion (E) 3.78 (0.11) 3.15 (0.28) 3.10 (0.37) 3.18 (0.90)
Agreeableness (A) 4.24 (0.13) 4.27 (0.32) 4.27 (0.18) 3.66 (0.72)
Neuroticism (N) 1.96 (0.20) 1.98 (0.39) 2.11 (0.21) 3.04 (0.88)
@@
Figure 1: Radar plot of the average Big Five personality scores for GPT-3.5-turbo, GPT-4o, GPT-5, and humans. Scores for the models are averaged over 20 repeated measurements.
Across all models, the 20 measurements yielded highly consistent scores, with relatively small standard deviations. Specifically, the standard deviations ranged from 0.11 to 0.20 for GPT-3.5-turbo, from 0.25 to 0.39 for GPT-4o, and from 0.16 to 0.37 for GPT-5. These values are substantially smaller than those observed in humans (0.66–0.90), indicating that each LLM exhibits a stable and internally consistent personality score. This consistency suggests that the measured personality traits reliably characterize each model.
In contrast, the human scores cluster around the mid-range of the scale, with larger standard deviations, reflecting substantial inter-individual variability in human personality traits. Compared to human averages, all LLMs exhibit higher scores in openness, conscientiousness, and agreeableness, as well as lower scores in neuroticism. This pattern suggests that LLMs tend to exhibit personality scores that are systematically shifted toward higher scores in certain traits compared to humans, potentially reflecting biases toward socially desirable response patterns present in their training data.
Next, we examine the patterns observed on individual personality dimensions. Overall, all models exhibited broadly similar tendencies, characterized by relatively high scores in openness, conscientiousness, and agreeableness, and low scores in neuroticism. This consistency across models suggests that contemporary LLMs share common personality scores as measured by the BFI framework. A particularly notable result is observed for conscientiousness. While GPT-3.5-turbo and GPT-4o show relatively similar mean scores for conscientiousness (4.06 and 4.12), GPT-5 exhibits a higher score (4.69). One possible explanation for this difference is the technical improvements reported for GPT-5, including a significant reduction in hallucinated responses and enhanced reasoning consistency compared to earlier models [Openai2025]. Such improvements may contribute to more reliable, goal-oriented, and internally consistent responses, which are reflected in higher conscientiousness scores.